# Disentangling the Halo: Joint Model for Measurements of the Kinetic Sunyaev-Zeldovich Effect and Galaxy-Galaxy Lensing

### v1

@@ 3.2 The Two Halo Term
where $A_{\rm 2h}$ is an additional nuisance parameter to account for any potential systematic modelling errors in the 2-halo term, $\xi^{\rm 2h}_{\rm hm}(r)$ is the 2-halo term of the halo-matter correlation function, related to the matter correlation function by linear halo bias $\xi^{\rm 2h}_{\rm hm}=b(M)\xi^{\rm 2h}_{\rm mm}$, and $\bar{\rho}_{\rm m}$ is the comoving average cosmic matter density. We expect $A_{\rm 2h}$ to be less than unity to match the low amplitude of the 2-halo GGL signal Leauthaud et al. (2017); Amon et al. (2023) compared to predictions from clustering in Planck $\Lambda$CDM. As can be seen in Figure 2, the fit to only the kSZ effect measurements prefers much larger values of $A_{\rm 2h}$ than unity which is not physically justifiable. We choose our upper prior bound for $A_{\rm 2h}$ to minimize the prior volume effect (e.g. Ried Guachalla et al., 2025b) and reduce potential bias on the physical parameters of interest. We note that $M$ denotes the mass of each halo. The correlation function is related to the power spectrum by the inverse Fourier transform
@@
[…]
where $P^{\rm lin}(k)$ is the linear matter power spectrum computed using the core cosmology library (pyccl) python package Chisari et al. (2019b), $dn/dM$ is the halo mass function Tinker et al. (2008), $b(M)$ is the linear halo bias Tinker et al. (2010), and $\hat{\rho}_{\rm NFW}(k)$ is the Fourier transform of the NFW profile. In this model we integrate over halo mass from $10^{10}M_{\odot}$ and $10^{15}M_{\odot}$. This completes the calculation of the 2-halo term of the total matter density profile. To compute the 2-halo contribution to the dark matter and baryon profiles we simply have
@@ 3.4 Modelling Galaxy-Galaxy Lensing
[…]
We compute the total matter excess surface mass density using the FFTLog algorithm implemented in pyccl Chisari et al. (2019b). With this result we have a complete consistent framework for modelling the both the kSZ effect and GGL measurements.
@@ 4 Results: joint model constraints
Figure 3: The baryon density profile (left), kSZ temperature profile at 150 GHz (top center) and 98 GHz (bottom center), and the GGL excess surface mass density profile (top right). We additionally show the fractional difference in $\Delta\Sigma$ relative to a model without any baryons (bottom right). The black points correspond to observations, the red corresponds to our fiducial model being fit to only the kSZ measurements, the blue corresponds to our fiducial model being fit to only the GGL measurements, and the purple corresponds to our joint fits to both the kSZ and GGL data. Dashed curves refer to the median (50th percentile) model of the posterior distribution and the bands denote the ±$\pm 1\sigma$ (32nd-68th percentiles) of the distribution of model curves. Lastly, solid curves denote the best-fit to the model. We show that the profile of the baryon density is tightly constrained by the joint fit relative to our kSZ or GGL only fits and qualitatively different compared to the data limited fits. We show that the joint fit of kSZ with GGL tightly constrains the baryon suppression in $\Delta\Sigma$ at small scales.
@@
We refer to $\vec{m}(\theta)$ as our model function for both $T_{\rm kSZ}$ and $\Delta\Sigma$ given our model parameters $\theta$. We define $\mathrm{Pr}(\theta)$ to be the priors on our model parameters $\theta$ which can be found in Table 1. We sample the posterior using an affine-invariant ensemble Markov Chain Monte-Carlo (MCMC) algorithm implemented in the package emcee Foreman-Mackey et al. (2013). To ensure convergence we ran the chains sufficiently long such that the Gelman-Rubin convergence parameter $\mathcal{R}$ was less than 1.01 for all model parameters Gelman & Rubin (1992).
@@ 4.1 The complementarity of kSZ and GGL
Figure 2 shows the posteriors for our model parameters obtained from fitting to the stacked kSZ profile data (red). The best-fit (minimum $\chi^{2}$) for this model is $\chi^{2}_{\rm min}/\mathrm{d.o.f}=21.8/14$ (PTE $=0.08$). As one would expect, the halo mass and amplitude of the 2-halo term are largely unconstrained when the model only has access to the kSZ profile. [^5] Note that we tested the impact of using a substantially flexible prior on the amplitude of the 2-halo term, $A_{\rm 2h}=\mathcal{U}(0.0,20.0)$. In this case, we do constrain the parameter, $A_{\rm 2h}=3.2^{+1.4}_{-2.4}$. For our main analysis, we keep the physically-motivated prior to avoid exploring large parts of unphysical parameter space. Complementary are the blue contours, which show the posterior from the GGL data. The GGL measurements offer a lot of constraining power on $M_{\rm halo}$ and $A_{\rm 2h}$ but don’t have much constraining power for $\beta$ or $x_{\rm c}$. This fit to the data gave a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=2.5/7$ (PTE $=0.92$) implying that the GGL measurements were over-fit by our model. Furthermore, the GGL measurements are a probe of the total matter distribution which is comprised of mostly dark matter making it difficult to infer substantial information of the baryon density profile.
The joint fit to the data is shown in purple. This fit yielded a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=24.8/25$ (PTE $=0.47$) implying a good fit to the data. In these purple contours we can see the overlapping regions of unconstrained red and blue contours leading to relatively tight constraints on all parameters except $x_{\rm c}$. If we fix $x_{\rm c}=0.5$, then we can see that $M_{\rm halo}$, $\beta$, and $A_{\rm 2h}$ are all tightly constrained (denoted by dashed black line in Figure 2) with little degeneracy. By jointly fitting to both kSZ and GGL data we have ‘disentangled’ baryons from dark matter.
@@ 4.2 Halo Mass
As a consequence of the suppression of $\Delta\Sigma$ caused by baryons, as seen in Figure 3 (lower right), neglecting to account for baryons can bias halo mass estimates $M_{\rm halo}$, compared to the true mass (e.g. Debackere et al., 2020; Cromer et al., 2022). This effect is demonstrated in Figure 4 where we show the posteriors of the halo mass for three analyses of the GGL data: fitting a model that neglects baryons ($f_{\rm b}=0\;\rightarrow\;f_{\rm cdm}=1$) to the GGL data, fitting our model which includes baryons to the same GGL data, and the same, where we jointly analyse GGL with kSZ profile measurements. We find that neglecting the impact of baryons lowers our estimate of the halo mass by $\sim 20\%$. Including the impact of baryons in the model but only using GGL data inflates our posterior, which is improved by the inclusion of the kSZ measurements. That is, a joint analysis of GGL and kSZ can account for the gas parameters, shrinking the width of the posterior back down, while the halo mass estimate undergoes a $\sim 20\%$ shift compared to the no-baryon model. We note that our joint kSZ+GGL constraints on $M_{\rm halo}$ are expected to be underestimated by $\lesssim 10\%$ compared to the mean halo mass of the sample $\langle M_{\rm halo}\rangle=3\times 10^{13}\;\rm M_{\odot}$ likely due to the exclusion of miscentering in our model (Skibba et al., 2011; Hoshino et al., 2015).
@@
A crucial step in using kSZ profile measurements as a benchmark for feedback models and simulations is an accurate constraint on the halo mass. The halo mass can change both the shape and the amplitude of the kSZ signal as seen in Figure 1. The ratio of kSZ signals modelled with halo masses of $10^{14}\;\rm M_{\odot}$ and $10^{13}\;\rm M_{\odot}$ is $\approx 2.1$ at $\theta=1^{\prime}$ and grows with increasing $\theta$, plateauing to $\approx 3.3$ at $\theta=6^{\prime}$. Assuming a certain halo mass estimated from stellar masses is potentially dangerous as stellar masses and SHMR can be biased (Yuan et al., 2021; Oyarzún et al., 2024). These kSZ profile measurements were jointly analysed with cosmic shear for the first time by Bigwood et al. (2024) to constrain the power spectrum and they found a best-fitting model with stronger baryonic feedback than hydrodynamical simulations predict. As discussed in Bigwood et al. (2024) (their Appendix B3), a major source of uncertainty in constraining the impact of baryonic feedback from kSZ and cosmic shear measurements, was the halo mass of the sample, which spans almost an order of magnitude: $M_{\rm halo}=0.8-30\times 10^{13}\;M_{\odot}$ (significantly larger than the range of masses plotted in Figure 4). For this reason, Bigwood et al. (2024) had to use a large prior for $M_{\rm halo}:\mathcal{U}[5\times 10^{12},7\times 10^{13}]$. Having such a large uncertainty on halo masses make it difficult to interpret the strength of baryonic feedback from the stacked kSZ profiles measured by Schaan et al. (2021); Hadzhiyska et al. (2024a) and future measurements to come. The methodology in this work can provide a more precise prior on the effective halo mass of the sample.
@@
We compare our ($2\sigma$) posteriors on the halo mass $\log_{10}(M_{\rm halo,200m}/\rm M_{\odot})=13.44\pm 0.12$ to the halo mass found by McCarthy et al. (2024) where they jointly analysed the kSZ and GGL measurements with the FLAMINGO simulation suite (Schaye et al., 2023). In this work we’ve used a mass definition of $M_{\rm 200m}$, to match McCarthy et al. (2024) we transform to a mass definition of $M_{\rm 500c}$ using an NFW profile (Navarro et al., 1997) and the concentration mass relation of Duffy et al. (2008). Our ($2\sigma$) posteriors on the halo mass are $\log_{10}(M_{\rm halo,500c}/\rm M_{\odot})=13.22\pm 0.12$ and the ($2\sigma$) posteriors on the halo mass from McCarthy et al. (2024) are $\log_{10}(M_{\rm 500c}/\rm M_{\odot})=13.34\pm 0.04$. We find that our estimate of the halo mass is $\sim 30\%$ lower than what was found by McCarthy et al. (2024).
Moreover, getting an unbiased estimate of the halo mass is crucial for modelling the tSZ effect (Moser et al., 2021, 2023) and the matter power spectrum suppression due to baryons (Debackere et al., 2020). Unlike the kSZ effect which scales linearly with halo mass $T_{\rm kSZ}(\theta)\propto M_{\rm halo}$, the tSZ effect has a much stronger scaling with halo mass $T_{\rm tSZ}(\theta)\propto M_{\rm halo}^{5/3}$. For this reason, an estimated halo mass that is biased high by $30\%$ will bias the model to over predict the amplitude of $T_{\rm tSZ}(\theta)$ by at least $50\%$. This potential bias arising from incorrect halo mass estimates could help alleviate discrepancies between the observed and predicted tSZ profiles in Moser et al. (2022) and tSZ clustering signals described in McCarthy et al. (2024). Similarly, a bias in halo mass estimates of $\sim 30\%$ could lead to an underestimation of the suppression of the matter power spectrum by $\sim 4\%$ at $k=1\;h/\mathrm{Mpc}$ (Debackere et al., 2020). By virtue of having an analytic model that can consistently model both the kSZ profile measurements and the GGL measurements we can completely side step any assumptions made to estimate the halo mass of the sample.
In future work we will improve our model by including satellites and accounting for miscentering which will yield an entirely unbiased estimate of the halo mass.
@@ 4.3 Baryonic Feedback
The kSZ effect probes the distribution of gas in the outskirts of halos. Only recently have direct measurements of the stacked kSZ effect profile had high enough SNR to be a powerful probe of the gas (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a). The ACT-BOSS kSZ measurements used in this analysis from Schaan et al. (2021) were first analysed with their corresponding tSZ measurements to study the thermodynamic properties of the halo by Amodeo et al. (2021) and have recently been jointly analysed with cosmic shear for the first time by Bigwood et al. (2024) to constrain the matter power spectrum. The joint analysis of Bigwood et al. (2024) found a best-fitting model that preferred a more extreme baryonic feedback scenario than hydrodynamical simulations predict. These same kSZ and GGL measurements were then analysed using the FLAMINGO simulations directly by McCarthy et al. (2024). Their analysis found that the kSZ measurements required more extreme baryonic feedback than the FLAMINGO simulations predicted by $7\sigma$. With this context in mind, we compare our joint kSZ+GGL constraints to the GNFW gas profile scaling relations derived from hydrodynamical simulations in Battaglia (2016).
@@
We show our ±$\pm 1\sigma$ estimates on $\beta$ and $M_{\rm halo}$ compared to 2 models from Battaglia (2016). The blue curve in Figure 5 is the Adiabatic model which corresponds to a simulation with only gravitational heating, the red curve is the AGN model which includes radiative cooling, star formation, supernova feedback, cosmic rays, and AGN feedback (more extreme feedback). A lower value of $\beta$ implies that the baryons are more spread out to larger radii implying stronger baryonic feedback. We find that our posteriors on halo mass and $\beta$ are consistent with the assumptions and findings of Amodeo et al. (2021) where they found $\beta=2.6^{+1.0}_{-0.6}$ (at fixed mean halo mass $\langle M_{\rm halo}\rangle=3\times 10^{13}\;\rm M_{\odot}$: light blue in Figure 5 with necessary correction discussed in Appendix B) when fitting a GNFW gas profile to the same kSZ measurements. [^6] We note that the modelling done in Amodeo et al. (2021) includes a mass-weighted average over the halo masses in their sample, for comparison we plot their constraints on $\beta$ at the mean halo mass in their sample in Figure 5. We find our constraints on $\beta$ are tighter than Amodeo et al. (2021) by a factor of 2 when fitting our model to the same data (kSZ Only fit in Table 1). This improvement in constraining $\beta$ is due to the way we normalize the baryon (gas) density profile (see Eq. 21) which effectively increases the importance of $\beta$ in the model. When we include GGL in the kSZ+GGL joint analysis our constraints on $\beta$ tighten further by 50% yielding a total factor of 3 improvement compared to Amodeo et al. (2021) while also giving a direct estimate on the halo mass $M_{\rm halo}$.
@@
[…]
@@
We note that our model neglects satellites, for this reason we are neglecting the potential systematic of ’miscentering’. We expect that not accounting for this effect causes an underestimation of the kSZ amplitude in simulations by up to $\sim 20\%$ for CMASS-like galaxies ($12.111.5<\log M_{*}<$ 12.1) as shown in Figure 6 of Hadzhiyska et al. (2023), at the current level of statistical power, this effect would not significantly change the conclusions of this work. We note that this effect is dependent on the satellite fraction of the sample which means that miscentering is more significant for lower mass galaxies ($11.510.9<\log M_{*}<$ 11.5) and less significant for more massive galaxies ($12.712.1<\log M_{*}<$ 12.7). Considering Figure 1, we can see that ignoring miscentering from satellites in our model would most likely mean our values for $\beta$ are overestimated by $\sim 0.2-0.4$ which would not significantly impact the conclusions from this study. There is of course, a possibility that miscentering could alter the shape of the kSZ profile further altering the value of $\beta$ changing our results. Including this effect in the model can become computationally expensive, but it has been modelled for observations of the tSZ effect (Popik et al., 2025). We leave this improvement to our model of kSZ+GGL as future work.
Figure 5: Scaling relations derived from Battaglia (2016) hydrodynamical simulations for the outer power law slope, $\beta$, of the baryon density profile in our model as a function of halo mass, $M_{\rm halo}$. The blue line corresponds to simulations with only gravitational heating and the red line corresponds to simulations with additional radiative cooling, star formation, supernova feedback, cosmic rays, and AGN feedback. Lower values of the outer power law slope $\beta$ indicate the baryons are more diffusely spread out in the density profile to the outermost regions of the halo. This suggests that the data prefers stronger baryonic feedback relative to the hydrodynamical simulations of Battaglia (2016). In our model we use a fiducial value for the baryon radius $r_{\rm b}=10\times r_{200}$ (purple), but we include an additional data point with $r_{\rm b}=20\times r_{200}$ to show the range of values $\beta$ could be with this model depending on the choice of $r_{\rm b}$.

### v2

@@ 3.2 The Two Halo Term
[…]
@@
$P^{\rm 2h}_{\rm mm}(k)=P^{\rm lin}(k)\left[\int dM^{\prime}\frac{dn}{dM^{\prime}}b(M^{\prime})\hat{\rho}_{\rm NFW}(k)\right]^{2}\;,$ (24)
where $P^{\rm lin}(k)$ is the linear matter power spectrum computed using the core cosmology library (pyccl) python package Chisari et al. (2019a), $dn/dM$ is the halo mass function Tinker et al. (2008), $b(M)$ is the linear halo bias Tinker et al. (2010), and $\hat{\rho}_{\rm NFW}(k)$ is the Fourier transform of the NFW profile. In this model we integrate over halo mass from $10^{10}M_{\odot}$ and $10^{15}M_{\odot}$. This completes the calculation of the 2-halo term of the total matter density profile. To compute the 2-halo contribution to the dark matter and baryon profiles we simply have
@@ 3.4 Modelling Galaxy-Galaxy Lensing
$\overline{\Sigma}(<R)=\frac{2}{R^{2}}\int dR^{\prime}\;R^{\prime}\;\Sigma(R^{\prime})\;.$ (36)
We compute the total matter excess surface mass density using the FFTLog algorithm implemented in pyccl Chisari et al. (2019a). With this result we have a complete consistent framework for modelling the both the kSZ effect and GGL measurements.
@@ 4 Results: joint model constraints
Figure 4: The baryon density profile (left), kSZ temperature profile at 150 GHz (top center) and 98 GHz (bottom center), and the GGL excess surface mass density profile (top right). We additionally show the fractional difference in $\Delta\Sigma$ relative to a model without any baryons (bottom right). The black points correspond to observations, the red corresponds to our fiducial model being fit to only the kSZ measurements, the blue corresponds to our fiducial model being fit to only the GGL measurements, and the purple corresponds to our joint fits to both the kSZ and GGL data. Dashed curves refer to the median (50th percentile) model of the posterior distribution and the bands denote the $\pm 1\sigma$ (32nd-68th percentiles) of the distribution of model curves. Lastly, solid curves denote the best-fit to the model. We show that the profile of the baryon density is tightly constrained by the joint fit relative to our kSZ or GGL only fits and qualitatively different compared to the data limited fits. We show that the joint fit of kSZ with GGL tightly constrains the baryon suppression in $\Delta\Sigma$ at small scales.
@@
We refer to $\vec{m}(\theta)$ as our model function for both $T_{\rm kSZ}$ and $\Delta\Sigma$ given our model parameters $\theta$. We define $\mathrm{Pr}(\theta)$ to be the priors on our model parameters $\theta$ which can be found in Table 1. We sample the posterior using an affine-invariant ensemble Markov Chain Monte-Carlo (MCMC) algorithm implemented in the package emcee Foreman-Mackey et al. (2013). To ensure convergence we ran the chains sufficiently long such that the Gelman-Rubin convergence parameter $\mathcal{R}$ was less than 1.01 for all model parameters Gelman and Rubin (1992).
@@ 4.1 The complementarity of kSZ and GGL
Figure 3 shows the posteriors for our model parameters obtained from fitting to the stacked kSZ profile data (red). The best-fit (minimum $\chi^{2}$) for this model is $\chi^{2}_{\rm min}/\mathrm{d.o.f}=21.8/16$ (PTE =0.15=0.15). As one would expect, the halo mass and amplitude of the 2-halo term are largely unconstrained when the model only has access to the kSZ profile. [^5] Note that we tested the impact of using a substantially flexible prior on the amplitude of the 2-halo term, $A_{\rm 2h}=\mathcal{U}(0.0,20.0)$. In this case, we do constrain the parameter, $A_{\rm 2h}=3.2^{+1.4}_{-2.4}$. For our main analysis, we keep the physically-motivated prior to avoid exploring large parts of unphysical parameter space. Complementary are the blue contours, which show the posterior from the GGL data. The GGL measurements offer a lot of constraining power on $M_{\rm halo}$ and $A_{\rm 2h}$ but don’t have much constraining power for $\beta$ or $x_{\rm c}$. This fit to the data gave a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=2.5/9$ (PTE =0.98=0.98) implying that the GGL measurements were over-fit by our model. Furthermore, the GGL measurements are a probe of the total matter distribution which is comprised of mostly dark matter making it difficult to infer substantial information of the baryon density profile.
The joint fit to the data is shown in purple. This fit yielded a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=24.8/26$ (PTE =0.53=0.53) implying a good fit to the data. In these purple contours we can see the overlapping regions of unconstrained red and blue contours leading to relatively tight constraints on all parameters except $x_{\rm c}$. If we fix $x_{\rm c}=0.5$, then we can see that $M_{\rm halo}$, $\beta$, and $A_{\rm 2h}$ are all tightly constrained (denoted by dashed black line in Figure 3) with little degeneracy. By jointly fitting to both kSZ and GGL data we have ‘disentangled’ baryons from dark matter.
@@ 4.2 Halo Mass
As a consequence of the suppression of $\Delta\Sigma$ caused by baryons, as seen in Figure 4 (lower right), neglecting to account for baryons can bias halo mass estimates $M_{\rm halo}$, compared to the true mass (e.g. Debackere et al., 2020; Cromer et al., 2022). This effect is demonstrated in Figure 5 where we show the posteriors of the halo mass for three analyses of the GGL data: fitting a model that neglects baryons ($f_{\rm b}=0\;\rightarrow\;f_{\rm cdm}=1$) to the GGL data, fitting our model which includes baryons to the same GGL data, and the same, where we jointly analyse GGL with kSZ profile measurements. We find that neglecting the impact of baryons lowers our estimate of the halo mass by $\sim 20\%$. Including the impact of baryons in the model but only using GGL data inflates our posterior, which is improved by the inclusion of the kSZ measurements. That is, a joint analysis of GGL and kSZ can account for the gas parameters, shrinking the width of the posterior back down, while the halo mass estimate undergoes a $\sim 20\%$ shift compared to the no-baryon model.
Our GGL+Baryons and kSZ+GGL constraints on $M_{\rm halo}$ are consistent with the mean halo mass $\langle M_{\rm halo}\rangle=3\times 10^{13}\;\rm M_{\odot}$ obtained with the SHMR Schaan et al. (2021) at the $1\sigma$ level but our GGL with no baryons constraint on halo mass shifts downward to agree at the $2\sigma$ level. We note that our mass estimate could be underestimated by $\lesssim 10\%$ due to the exclusion of miscentering in our model (Skibba et al., 2011; Hoshino et al., 2015). In future work we will improve our model by including satellites and accounting for miscentering which will yield an unbiased estimate of the halo mass.
A crucial step in using kSZ profile measurements as a benchmark for feedback models and simulations is an accurate constraint on the halo mass. The halo mass can change both the shape and the amplitude of the kSZ signal as seen in Figure 2. The ratio of kSZ signals modelled with halo masses of $10^{14}\;\rm M_{\odot}$ and $10^{13}\;\rm M_{\odot}$ is $\approx 2.1$ at $\theta=1^{\prime}$ and grows with increasing $\theta$, plateauing to $\approx 3.3$ at $\theta=6^{\prime}$. Assuming a certain halo mass estimated from stellar masses is potentially dangerous as stellar masses and SHMR can be biased (Yuan et al., 2021; Oyarzún et al., 2024).
@@
These kSZ profile measurements were jointly analysed with cosmic shear for the first time by Bigwood et al. (2024) to constrain the power spectrum and they found a best-fitting model with stronger baryonic feedback than hydrodynamical simulations predict. As discussed in Bigwood et al. (2024) (their Appendix B3), a major source of uncertainty in constraining the impact of baryonic feedback from kSZ and cosmic shear measurements, was the halo mass of the sample, which spans almost an order of magnitude: $M_{\rm halo}=0.8-30\times 10^{13}\;M_{\odot}$ (significantly larger than the range of masses plotted in Figure 5). For this reason, Bigwood et al. (2024) had to use a large prior for $M_{\rm halo}:\mathcal{U}[5\times 10^{12},7\times 10^{13}]$. Having such a large uncertainty on halo masses makes it difficult to interpret the strength of baryonic feedback from the stacked kSZ profiles measured by Schaan et al. (2021); Hadzhiyska et al. (2024a) and future measurements to come. The methodology in this work can provide a more precise prior on the effective halo mass of the sample.
@@
We compare our ($2\sigma$) posteriors on the halo mass $\log_{10}(M_{\rm halo,200m}/\rm M_{\odot})=13.44\pm 0.12$ to the halo mass found by McCarthy et al. (2024) where they jointly analysed the kSZ and GGL measurements with the FLAMINGO simulation suite (Schaye et al., 2023). In this work we’ve used a mass definition of $M_{\rm 200m}$, to match McCarthy et al. (2024) we transform to a mass definition of $M_{\rm 500c}$ using an NFW profile (Navarro et al., 1997) and the concentration mass relation of Duffy et al. (2008). Our ($2\sigma$) posteriors on the halo mass are $\log_{10}(M_{\rm halo,500c}/\rm M_{\odot})=13.22\pm 0.12$ and the ($2\sigma$) posteriors on the halo mass from McCarthy et al. (2024) are $\log_{10}(M_{\rm 500c}/\rm M_{\odot})=13.34\pm 0.04$. We find that our estimate of the halo mass is slightly lower than what was found by McCarthy et al. (2024) but still statistically consistent at the $2\sigma$ level. We note that there is additional uncertainty introduced by our method of converting between mass definitions due to the fact that the baryons are not described by an NFW profile, but the baryon fraction at $r_{\rm 500c}$ is lower than the cosmic baryon fraction and should only introduce uncertainty at the few percent level.
@@ 4.3 Baryonic Feedback
The kSZ effect probes the distribution of gas in the outskirts of halos. Only recently have direct measurements of the stacked kSZ effect profile had high enough SNR to be a powerful probe of the gas (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b). The ACT-BOSS kSZ measurements used in this analysis from Schaan et al. (2021) were first analysed with their corresponding tSZ measurements to study the thermodynamic properties of the halo by Amodeo et al. (2021) and have recently been jointly analysed with cosmic shear for the first time by Bigwood et al. (2024) to constrain the matter power spectrum. The joint analysis of Bigwood et al. (2024) found a best-fitting model that preferred a more extreme baryonic feedback scenario than hydrodynamical simulations predict. These same kSZ and GGL measurements were then analysed using the FLAMINGO simulations directly by McCarthy et al. (2024). Their analysis found that the kSZ measurements required more extreme baryonic feedback than the FLAMINGO simulations predicted by $7\sigma$. With this context in mind, we compare our joint kSZ+GGL constraints to the GNFW gas profile scaling relations derived from hydrodynamical simulations in Battaglia (2016).
@@
We show our $\pm 1\sigma$ estimates on $\beta$ and $M_{\rm halo}$ compared to 2 models from Battaglia (2016). The blue curve in Figure 6 is the Adiabatic model which corresponds to a simulation with only gravitational heating, the red curve is the AGN model which includes radiative cooling, star formation, supernova feedback, cosmic rays, and AGN feedback (more extreme feedback). A lower value of $\beta$ implies that the baryons are more spread out to larger radii implying stronger baryonic feedback. We find that our posteriors on halo mass and $\beta$ are consistent with the assumptions and findings of Amodeo et al. (2021) where they found $\beta=2.6^{+1.0}_{-0.6}$ (at fixed mean halo mass $\langle M_{\rm halo}\rangle=3\times 10^{13}\;\rm M_{\odot}$: light blue in Figure 6 with necessary correction discussed in Appendix B) when fitting a GNFW gas profile to the same kSZ measurements. [^6] We note that the modelling done in Amodeo et al. (2021) includes a mass-weighted average over the halo masses in their sample, for comparison we plot their constraints on $\beta$ at the mean halo mass in their sample in Figure 6. We find our constraints on $\beta$ are tighter than Amodeo et al. (2021) by a factor of 2 when fitting our model to the same data (kSZ Only fit in Table 1). This improvement in constraining $\beta$ is due to the way we normalize the baryon (gas) density profile (see Eq. 21) which effectively increases the importance of $\beta$ in the model. When we include GGL in the kSZ+GGL joint analysis our constraints on $\beta$ tighten further by 50% yielding a total factor of 3 improvement compared to Amodeo et al. (2021) while also giving a direct estimate on the halo mass $M_{\rm halo}$.
@@
The difference in the suppression of $\Delta\Sigma$ between the older Illustris simulation and the TNG-300 simulation can be explained by the differences in subgrid model implementations for feedback. The Illustris model is known to have more extreme feedback than the TNG-300 model but is also known to have shortcomings in predicting observations such as gas fractions, galaxy color and morphologies, and several cluster properties (Nelson et al., 2019; Kauffmann et al., 2019; Rodriguez-Gomez et al., 2019). Our kSZ+GGL best-fit suggests a suppression of $\Delta\Sigma$ that gradually increases from 0% at 10 Mpc/$h$ to $15\%$ for $R\lesssim 1\;\mathrm{Mpc}/h$ (in comoving units) which appears to more closely reflect the Illustris feedback model than the IllustrisTNG-300 feedback model. Newer measurements of the kSZ effect profile in DESI-Y1 data also appear to have a shape that agrees with the older Illustris feedback model more than the IllustrisTNG-300 feedback model (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b).
@@
We note that our model neglects satellites, for this reason we are neglecting the potential systematic of ’miscentering’. We expect that not accounting for this effect causes an underestimation of the kSZ amplitude in simulations by up to $\sim 20\%$ for CMASS-like galaxies ($11.5<\log M_{*}<12.1$) as shown in Figure 6 of Hadzhiyska et al. (2023), at the current level of statistical power, this effect would not significantly change the conclusions of this work. We note that this effect is dependent on the satellite fraction of the sample which means that miscentering is more significant for lower mass galaxies ($10.9<\log M_{*}<11.5$) and less significant for more massive galaxies ($12.1<\log M_{*}<12.7$). Considering Figure 2, we can see that ignoring miscentering from satellites in our model would most likely mean our values for $\beta$ are overestimated by $\sim 0.2-0.4$ which would not significantly impact the conclusions from this study. There is of course, a possibility that miscentering could alter the shape of the kSZ profile further altering the value of $\beta$ changing our results. Including this effect in the model can become computationally expensive, but it has been modelled for observations of the tSZ effect (Popik et al., 2025). We leave this improvement to our model of kSZ+GGL as future work.
Figure 6: Scaling relations derived from Battaglia (2016) hydrodynamical simulations for the outer power law slope, $\beta$, of the baryon density profile in our model as a function of halo mass, $M_{\rm halo}$. The blue line corresponds to simulations with only gravitational heating and the red line corresponds to simulations with additional radiative cooling, star formation, supernova feedback, cosmic rays, and AGN feedback. Lower values of the outer power law slope $\beta$ indicate the baryons are more diffusely spread out in the density profile to the outermost regions of the halo. This suggests that the data prefers stronger baryonic feedback relative to the hydrodynamical simulations of Battaglia (2016). In our model we use a fiducial value for the baryon radius $r_{\rm b}=10\times r_{200}$ (purple), but we include an additional data point with $r_{\rm b}=20\times r_{200}$ to show the range of values $\beta$ could be with this model depending on the choice of $r_{\rm b}$. All fits plotted from this work allow $x_{\rm c}$ to be a free parameter.